OpenStudy (anonymous):

Anyone know what one is an example of an irrational number? •negative square root of 17 •square root of 16 •negative square root of 9 •square root of 1

OpenStudy (study312):

Irrational is infinite like pi

OpenStudy (dayakar):

do know which are irrational numbers

OpenStudy (anonymous):

No

OpenStudy (dayakar):

can u divide 1/4 7/5

OpenStudy (anonymous):

5/28

OpenStudy (dayakar):

1/4= 7/5= each one seperately

OpenStudy (anonymous):

0.25 and 1.4

OpenStudy (dayakar):

very good these are called terminated decimals 1/3= can u find it

OpenStudy (anonymous):

1.3 repeating

OpenStudy (neonumbrella5115):

A quick response is anything impossible or imaginary, can be referred to irrational

OpenStudy (dayakar):

very good any number which is not terminating and not repeating is called irrational numer

OpenStudy (dayakar):

example: \[\sqrt{2 }, \sqrt{3}, \sqrt{11}, \pi...etc\]

OpenStudy (anonymous):

Thank you!

OpenStudy (neonumbrella5115):

Does that help you choose an answer?

OpenStudy (dayakar):

\[\sqrt{1}\] = ?

OpenStudy (anonymous):

I don't know what the square root of 1 is and yes it kind of helps me choose an answer But I'm still so confused

OpenStudy (neonumbrella5115):

The square root on 1 is one because one times one is one. ;)

OpenStudy (anonymous):

Oh haha! Thank you

OpenStudy (wolf1728):

Why not look up the definition? (of a number) not capable of being expressed exactly as a ratio of two integers. So, numbers like pi, 'e', square root of two, etc are irrational.